OpenStudy (anonymous):

Lim X-> (infinity): x/sqrt(x^2-1) ...having trouble working out the square root on the bottom

OpenStudy (anonymous):

tried to divide by highest degree of x in the denominator by pulling it from the square root making it X^4 unfortunately this makes numerator approach 0 ...which is not correct I don't think

OpenStudy (anonymous):

try making the radical an exponent and move it to the numerator

OpenStudy (anonymous):

forget the \(-1\) part, it is totally unimportant and just write \[\frac{x}{\sqrt{x^2}}=\frac{x}{x}=1\]

OpenStudy (anonymous):

I guess thats true if you are approaching infinity